# Stability analysis through folds: An end-loaded elastica with a lever arm

### v5

@@ 4 Numerical Examples
The distinguished bifurcation diagrams facilitate the determination of the index change at a fold in the varying parameter. If there are no folds, the stability remains the same for all solutions along the branch. At folds, the stability transitions usually occur, and they were evaluated by examining the manner in which the fold occurs, as depicted in Figure 1. We use the straight trivial solutions from the previous section as initial solutions to generate the family of equilibria by varying parameters. Subsequently, we assess the stability of the resulting equilibria. Numerical continuation is performed using AUTO-07p [39], which employs pseudo-arclength continuation [1], and is capable of detecting singularities like folds and bifurcations as the parameter is varied. In these examples, the analysis is restricted solely to simple folds.
@@ 4.1.1 Rotating Arm
In the last step, the boundary condition (38b) is applied. This ordinate $\mathbf{m}(l).\mathbf{e}_{z}$ is evaluated from the numerical continuation solutions and plotted against the parameter $\psi$ to generate the associated distinguished bifurcation diagram, as shown in Figure 4. This plot exhibits folds, indicating stability transitions. The direction of change in index near folds is inferred from Figure 1(b) and is indicated in Figures 4. The equilibrium at $\psi=\pi$ is equivalent to the case of $\epsilon<0$ in Section 3.2, where $P<P_{cr,1}^{-}$ corresponds to index zero and agrees with the diagrams. The dynamical aspects of the snapping motion are not discussed in this paper, and for further details, refer to [40, 18]. The solutions and their projections exhibit $2\pi$- periodic characteristics. Figure 4(a) also displays the elastica configurations at intermediate values of $\psi$. The potential snapping motion depicting the initial and final configurations is also shown. This motion depends on the current state and the direction of current parameter change, and the snap jump is indicated by a red arrow.
@@ 5 Conclusion
The theory of distinguished bifurcation diagrams was extended to problems characterized by fixed-free ends for their stability analysis. Using a combination of bifurcation diagrams and fold information, the stability of elastica subjected to an end load through a rigid lever arm was studied. The stability is determined through a qualitative examination of the plots, without performing a rigorous analysis such as conjugate point computations or eigenvalue determination. Nevertheless, the stability index of at least one solution along a branch must be established, which was determined using bifurcation analysis. Despite their simplicity and relative abundance, elastica systems with a lever arm have received limited attention. We presented several instances of multi-stability arising in these special elasticas by tuning various parameters. Bi-stability or multi-stability is extensively investigated for engineering applications such as microfluidic devices [41], soft robots [42, 43] and MEMS devices [44, 45]. Many modern functional materials [46] have been developed, where electric, magnetic, optical, thermal, or solvent-based stimuli can be used to tune system parameters near instability, and thereby enabling actuation. The presented elastica system holds significant potential for designing variety of innovative mechanisms. The distinguished bifurcation diagrams can serve as an invaluable tool in their design. A natural extension of this work would be to apply a similar framework to three-dimensional elastic rods using the Kirchhoff rod theory [47].
@@
We restricted our analysis to parameters that produce equilibria with a maximum index of $2$. Future work could explore parameters that generate higher-index equilibria. For example, larger values of $P$ would generate equilibria with higher indices. Another key aspect of the current findings is the system’s periodic nature with respect to the rotation parameters. In contrast to problems with fixed-fixed ends in [9], the present system exhibited periodicity with respect to the rotation parameters when one of its fixed ends is set free. The free end allows for more freedom, limiting the possibility of configurations with higher stored elastic energy. But, increasing the parameters, such as $P$ allows their production.
@@
In this study, we presented the equilibria resulting from the quasi-static control of parameters and drew conclusions on the equilibria after snap-back instability without detailing the transition path. In some instances, multiple stable equilibria are possible after the snap-back instability, and exact equilibrium to which the system transitions cannot be concluded. In this case, the study of dynamical aspects of snapping would be more advantageous [18, 40]. The most likely equilibrium depends on the energy released during snapping as well as the system’s dissipation. Incorporating isoperimetric constraints, where the position of the free end is fixed, while it is free to rotate, would lead to interesting scenarios. It has already been proved that these constraints would have no effect on the ordinate of bifurcation diagrams for the case of fixed-fixed ends [9], and and it remains to be verified whether the same holds in the present case. The current analysis focuses solely on simple folds. An extension to degenerate scenarios involving non-simple folds would be a valuable direction for future work.

### v6

@@ 4 Numerical Examples
The distinguished bifurcation diagrams facilitate the determination of the index change at a fold in the varying parameter. If there are no folds, the stability remains the same for all solutions along the branch. At folds, the stability transitions usually occur, and they were evaluated by examining the manner in which the fold occurs, as depicted in Figure 1. We use the straight trivial solutions from the previous section as initial solutions to generate the family of equilibria by varying parameters. Subsequently, we assess the stability of the resulting equilibria. Numerical continuation is performed using AUTO-07p [doedel2007auto], which employs pseudo-arclength continuation [DOEDEL1991a], and is capable of detecting singularities like folds and bifurcations as the parameter is varied. In these examples, the analysis is restricted solely to simple folds.
@@ 4.1.1 Rotating Arm
In the last step, the boundary condition (38b) is applied. This ordinate $\mathbf{m}(l).\mathbf{e}_{z}$ is evaluated from the numerical continuation solutions and plotted against the parameter $\psi$ to generate the associated distinguished bifurcation diagram, as shown in Figure 4. This plot exhibits folds, indicating stability transitions. The direction of change in index near folds is inferred from Figure 1(b) and is indicated in Figures 4. The equilibrium at $\psi=\pi$ is equivalent to the case of $\epsilon<0$ in Section 3.2, where $P<P_{cr,1}^{-}$ corresponds to index zero and agrees with the diagrams. The dynamical aspects of the snapping motion are not discussed in this paper, and for further details, refer to [snyder1990dynamics, Armanini2017]. The solutions and their projections exhibit $2\pi$- periodic characteristics. Figure 4(a) also displays the elastica configurations at intermediate values of $\psi$. The potential snapping motion depicting the initial and final configurations is also shown. This motion depends on the current state and the direction of current parameter change, and the snap jump is indicated by a red arrow.
@@ 5 Conclusion
The theory of distinguished bifurcation diagrams was extended to problems characterized by fixed-free ends for their stability analysis. Using a combination of bifurcation diagrams and fold information, the stability of elastica subjected to an end load through a rigid lever arm was studied. The stability is determined through a qualitative examination of the plots, without performing a rigorous analysis such as conjugate point computations or eigenvalue determination. Nevertheless, the stability index of at least one solution along a branch must be established, which was determined using bifurcation analysis. Despite their simplicity and relative abundance, elastica systems with a lever arm have received limited attention. We presented several instances of multi-stability arising in these special elasticas by tuning various parameters. Bi-stability or multi-stability is extensively investigated for engineering applications such as microfluidic devices [Hilber2016], soft robots [Laschi2012, MajidiCarmel2014] and MEMS devices [zhang2007snap, maurini2007distributed]. Many modern functional materials [shen2020stimuli] have been developed, where electric, magnetic, optical, thermal, or solvent-based stimuli can be used to tune system parameters near instability, and thereby enabling actuation. The presented elastica system holds significant potential for designing variety of innovative mechanisms. The distinguished bifurcation diagrams can serve as an invaluable tool in their design. A natural extension of this work would be to apply a similar framework to three-dimensional elastic rods using the Kirchhoff rod theory [Dichmann1996].
@@
We restricted our analysis to parameters that produce equilibria with a maximum index of $2$. Future work could explore parameters that generate higher-index equilibria. For example, larger values of $P$ would generate equilibria with higher indices. Another key aspect of the current findings is the system’s periodic nature with respect to the rotation parameters. In contrast to problems with fixed-fixed ends in [Hoffman2005], the present system exhibited periodicity with respect to the rotation parameters when one of its fixed ends is set free. The free end allows for more freedom, limiting the possibility of configurations with higher stored elastic energy. But, increasing the parameters, such as $P$ allows their production.
@@
In this study, we presented the equilibria resulting from the quasi-static control of parameters and drew conclusions on the equilibria after snap-back instability without detailing the transition path. In some instances, multiple stable equilibria are possible after the snap-back instability, and exact equilibrium to which the system transitions cannot be concluded. In this case, the study of dynamical aspects of snapping would be more advantageous [Armanini2017, snyder1990dynamics]. The most likely equilibrium depends on the energy released during snapping as well as the system’s dissipation. Incorporating isoperimetric constraints, where the position of the free end is fixed, while it is free to rotate, would lead to interesting scenarios. It has already been proved that these constraints would have no effect on the ordinate of bifurcation diagrams for the case of fixed-fixed ends [Hoffman2005], and it remains to be verified whether the same holds in the present case. The current analysis focuses solely on simple folds. An extension to degenerate scenarios involving non-simple folds would be a valuable direction for future work.